### v1

@@ B.2 Additional Results
  • All specifications include user fixed effects. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of observations is in parentheses after each sample definition. Robust standard errors in parentheses are clustered at the user level.
Table B.3: Effect of ChatGPT Ban on GitHub Output — March 27–30 (Pre) – April 3–6 (Post)
Output Quantity Output Quality Task Complexity
Output Output (Shen, 2023) Commit Issue closed Bug fix ratio PR merged PR merge ratio Files edited per PR Lines added per PR
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Treated $\times$ Post 0.0110 0.0128 0.0111 -0.0005 0.0005 0.0059 0.0065 0.0103 0.0251
(0.0042) (0.0042) (0.0042) (0.0011) (0.0004) (0.0015) (0.0017) (0.0034) (0.0075)
Post -0.0149 -0.0153 -0.0140 -0.0003 $0$0.0030 -0.0031 -0.0056 -0.0119
(0.0019) (0.0019) (0.0018) (0.0005) (0.0002) (0.0008) (0.0009) (0.0017) (0.0039)
Overall (N = 290,992) Dep. var. mean 0.2286 0.2283 0.2215 0.0142 0.0026 0.0371 0.0396 0.0698 0.1593
Treated $\times$ Post 0.0181 0.0176 0.0168 0.0027 0.0008 0.0099 0.0109 0.0176 0.0415
(0.0061) (0.0061) (0.0060) (0.0013) (0.0004) (0.0018) (0.0020) (0.0044) (0.0095)
Post -0.0218 -0.0219 -0.0212 -0.0004 -0.0003 -0.0042 -0.0042 -0.0086 -0.0201
(0.0027) (0.0027) (0.0027) (0.0005) (0.0002) (0.0010) (0.0011) (0.0023) (0.0052)
Less experienced (N = 149,728) Dep. var. mean 0.2319 0.2316 0.2273 0.0089 0.0013 0.0265 0.0281 0.0549 0.1259
Treated $\times$ Post 0.0036 0.0079 0.0053 -0.0043 0.0001 0.0013 0.0015 0.0022 0.0070
(0.0059) (0.0059) (0.0058) (0.0020) (0.0008) (0.0026) (0.0029) (0.0053) (0.0118)
Post -0.0078 -0.0083 -0.0066 -0.0001 0.0003 -0.0018 -0.0021 -0.0024 -0.0035
(0.0025) (0.0025) (0.0024) (0.0008) (0.0003) (0.0012) (0.0014) (0.0026) (0.0057)
Experienced (N = 141,264) Dep. var. mean 0.2251 0.2249 0.2155 0.0198 0.0039 0.0483 0.0518 0.0857 0.1947
Treated $\times$ Post 0.0147 0.0151 0.0188 -0.0095 -0.0036 0.0002 0.0011 0.0008 -0.0030
(0.0109) (0.0108) (0.0106) (0.0040) (0.0017) (0.0050) (0.0058) (0.0106) (0.0233)
Post -0.0152 -0.0149 -0.0145 -0.0022 0.0004 -0.0023 -0.0027 -0.0040 -0.0059
(0.0046) (0.0046) (0.0045) (0.0016) (0.0007) (0.0024) (0.0029) (0.0047) (0.0107)
Pkg. contributor (N = 47,360) Dep. var. mean 0.2644 0.2645 0.2522 0.0267 0.0052 0.0635 0.0684 0.1087 0.2480
@@
  • All specifications include user fixed effects. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of observations is in parentheses after each sample definition. Robust standard errors in parentheses are clustered at the user level.
Table B.4: Alternative Outcomes
Output Quantity Task Choice Placebo
Event Output (Holub and Thies, 2023) Pull request (PR) Release Interactive activity Easy issue ratio Wiki page Public repo
(1) (2) (3) (4) (5) (6) (7) (8)
Pre Mar 27–28 – Post Apr 3–4
Overall (N = 145,496) Treated $\times$ Post 0.0224 0.0189 0.0066 -0.0017 0.0023 $0$0.0001 -0.0004
(0.0065) (0.0058) (0.0028) (0.0011) (0.0029) (0.0003) (0.0004) (0.0010)
Less experienced (N = 74,864) Treated $\times$ Post 0.0313 0.0257 0.0119 -0.0003 0.0025 -0.0002 -0.0006 0.0001
(0.0091) (0.0081) (0.0032) (0.0013) (0.0032) (0.0003) (0.0006) (0.0016)
Experienced (N = 70,632) Treated $\times$ Post 0.0134 0.0119 0.0009 -0.0032 0.0019 0.0002 0.0006 -0.0011
(0.0094) (0.0083) (0.0046) (0.0017) (0.0051) (0.0006) (0.0006) (0.0010)
Pkg. contributor (N = 23,680) Treated $\times$ Post 0.0363 0.0274 -0.0003 -0.0037 -0.0025 -0.0018 -0.0004 0.0005
(0.0166) (0.0154) (0.0094) (0.0033) (0.0105) (0.0009) (0.0013) (0.0020)
Pre Mar 27–30 – Post Apr 3–6
Overall (N = 290,992) Treated $\times$ Post 0.0190 0.0147 0.0061 -0.0005 0.0011 0.0002 0.0001 0.0002
(0.0050) (0.0045) (0.0020) (0.0008) (0.0022) (0.0002) (0.0003) (0.0007)
Less experienced (N = 149,728) Treated $\times$ Post 0.0244 0.0186 0.0107 -0.0005 0.0016 0.0001 -0.0004 0.0004
(0.0069) (0.0062) (0.0024) (0.0011) (0.0024) (0.0002) (0.0004) (0.0011)
Experienced (N = 141,264) Treated $\times$ Post 0.0136 0.0109 0.0011 -0.0006 0.0004 0.0003 0.0006 -0.0001
(0.0072) (0.0064) (0.0033) (0.0013) (0.0039) (0.0004) (0.0005) (0.0007)
Pkg. contributor (N = 47,360) Treated $\times$ Post 0.0337 0.0195 -0.0034 0.0009 0.0005 -0.0009 -0.0007 0.0001
(0.0125) (0.0115) (0.0066) (0.0024) (0.0081) (0.0005) (0.0011) (0.0011)
@@
  • All specifications include user fixed effects. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of observations is in parentheses after each sample definition. Robust standard errors in parentheses are clustered at the user level.
Table B.5: Event Counts
Output Quantity Output Quality
Ln(1 +Events) Ln(1 + Output) Ln(1 + Output (Holub and Thies, 2023)) Ln(1 + Output (Shen, 2023)) Ln(1 + Commits) Ln(1 + Issues closed) Ln(1 + PR merged)
(1) (2) (3) (4) (5) (6) (7)
Pre Mar 27–28 – Post Apr 3–4
Overall (N = 145,496) Treated $\times$ Post 0.0247 0.0178 0.0203 0.0244 0.0166 0.0001 0.0069
(0.0097) (0.0104) (0.0104) (0.0107) (0.0102) (0.0013) (0.0021)
Less experienced (N = 74,864) Treated $\times$ Post 0.0412 0.0368 0.0389 0.0420 0.0337 0.0010 0.0105
(0.0134) (0.0139) (0.0140) (0.0142) (0.0137) (0.0014) (0.0024)
Experienced (N = 70,632) Treated $\times$ Post 0.0070 -0.0033 -0.0001 0.0050 -0.0022 -0.0008 0.0027
(0.0141) (0.0156) (0.0156) (0.0161) (0.0153) (0.0023) (0.0035)
Pkg. contributor (N = 23,680) Treated $\times$ Post 0.0350 0.0138 0.0149 0.0247 0.0159 -0.0044 0.0037
(0.0268) (0.0313) (0.0313) (0.0322) (0.0307) (0.0045) (0.0074)
Pre Mar 27–30 – Post Apr 3–6
Overall (N = 290,992) Treated $\times$ Post 0.0264 0.0212 0.0237 0.0271 0.0206 -0.0001 0.0064
(0.0076) (0.0080) (0.0081) (0.0083) (0.0079) (0.0010) (0.0016)
Less experienced (N = 149,728) Treated $\times$ Post 0.0366 0.0345 0.0353 0.0386 0.0316 0.0024 0.0102
(0.0105) (0.0109) (0.0110) (0.0112) (0.0108) (0.0012) (0.0018)
Experienced (N = 141,264) Treated $\times$ Post 0.0157 0.0067 0.0110 0.0146 0.0087 -0.0031 0.0021
(0.0110) (0.0118) (0.0118) (0.0123) (0.0116) (0.0018) (0.0026)
Pkg. contributor (N = 47,360) Treated $\times$ Post 0.0482 0.0205 0.0214 0.0335 0.0225 -0.0053 0.0024
(0.0207) (0.0234) (0.0234) (0.0244) (0.0230) (0.0035) (0.0049)
@@
  • All specifications include user fixed effects. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of observations is in parentheses after each sample definition. Robust standard errors in parentheses are clustered at the user level.
Figure B.2: HonestDiD – Relative Magnitude Bounds
Notes: This figure displays HonestDiD robust confidence interval sets of the average treatment effect across the first two post-treatment periods for different relative magnitude bounds $\Delta^{RM}\left(\bar{M}\right)$, as suggested by Rambachan and Roth (2023). Panel A displays confidence interval sets for “less experienced” GitHub user accounts and Panel B for (i) “experienced” accounts and (ii) accounts that are the owner and/or contributor to a programming package repository (“Pkg. contributor”); 95% (90%) robust confidence intervals are depicted in light (dark) grey for relative magnitude bounds and in light (dark) red for original treatment effect estimates.
Appendix C GitHub Repository–User-Level Data
C.1 Descriptive Statistics
Table C.1: Descriptive Statistics

### v2

@@ C.2 Additional Results
  • All specifications include user–fixed effects and day-of-the-week–fixed effects. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0062 0.0102 0.0090 -0.0045** 0.0045
(N = 290,864) (0.0076) (0.0075) (0.0080) (0.0022) (0.0030)
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced Treated $\times$ Post 0.0193* 0.0216** 0.0203* -0.0017 0.0118***
(N = 149,680) (0.0107) (0.0107) (0.0111) (0.0023) (0.0035)
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced Treated $\times$ Post -0.0079 -0.0020 -0.0030 -0.0077* -0.0034
(N = 141,184) (0.0107) (0.0106) (0.0116) (0.0039) (0.0052)
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor Treated $\times$ Post 0.0181 0.0164 0.0270 -0.0151* 0.0113
(N = 47,328) (0.0205) (0.0200) (0.0215) (0.0084) (0.0104)
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post $0$0.0039 0.0220 -0.0004 0.0006
(N = 290,864) (0.0034) (0.0148) (0.0135) (0.0004) (0.0043)
Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced Treated $\times$ Post 0.0040 0.0211 0.0456*** -0.0004 0.0037
(N = 149,680) (0.0040) (0.0186) (0.0164) (0.0004) (0.0045)
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced Treated $\times$ Post -0.0046 -0.0324 -0.0031 -0.0003 -0.0028
(N = 141,184) (0.0056) (0.0237) (0.0221) (0.0008) (0.0077)
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor Treated $\times$ Post -0.0221* -0.0933* 0.0226 -0.0021* 0.0127
(N = 47,328) (0.0117) (0.0477) (0.0422) (0.0011) (0.0158)
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393
@@
  • All specifications include user–fixed effects, day-of-the-week–fixed effects and country-specific linear time trends. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Notes: Event-study estimates across outcomes for “experienced” GitHub user accounts (created before 2017). The sample period spans March 27–30 (Pre) and April 3–6 (Post). All specifications include user, time, and day-of-the-week fixed effects. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). 95% (90%) confidence intervals for robust standard errors clustered at the user level are depicted in light (dark) grey.
Notes: Event-study estimates across outcomes for “pkg. contributor” GitHub user accounts (owner and/or contributor to a programming package repository). The sample period spans March 27–30 (Pre) and April 3–6 (Post). All specifications include user, time, and day-of-the-week fixed effects. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). 95% (90%) confidence intervals for robust standard errors clustered at the user level are depicted in light (dark) grey.
(1) (2) (3) (4) (5)
A: Additional Outcomes
Any Event Commit PR merged PR merge ratio (Holub & Thies, 2023) Avg. files edited per merged PR
Overall Treated $\times$ Post 0.0091 0.0084 0.0047 0.0017 0.0119**
(N = 290,864) (0.0090) (0.0075) (0.0031) (0.0018) (0.0060)
Dep. var. mean 0.3358 0.2215 0.0371 0.0130 0.0630
Less experienced Treated $\times$ Post 0.0282** 0.0220** 0.0118*** 0.0037 0.0224***
(N = 149,680) (0.0124) (0.0106) (0.0035) (0.0024) (0.0074)
Dep. var. mean 0.3191 0.2272 0.0265 0.0122 0.0498
Experienced Treated $\times$ Post -0.0116 -0.0061 -0.0029 -0.0004 0.0005
(N = 141,184) (0.0133) (0.0105) (0.0053) (0.0028) (0.0098)
Dep. var. mean 0.3534 0.2155 0.0483 0.0138 0.0769
Pkg. contributor Treated $\times$ Post 0.0311 0.0242 0.0107 -0.0024 0.0277
(N = 47,328) (0.0233) (0.0199) (0.0107) (0.0050) (0.0195)
Dep. var. mean 0.4088 0.2521 0.0635 0.0171 0.0971
B: Intensive Margin
Outputs Outputs (Shen, 2023) Outputs (Holub & Thies, 2023) PRs merged PRs opened
Overall Treated $\times$ Post 0.0122 0.0164 0.0180 0.0053* 0.0010
(N = 290,864) (0.0145) (0.0146) (0.0149) (0.0032) (0.0030)
Dep. var. mean 0.3706 0.3727 0.4215 0.0347 0.0345
Less experienced Treated $\times$ Post 0.0407** 0.0464** 0.0478** 0.0144*** 0.0041
(N = 149,680) (0.0190) (0.0191) (0.0193) (0.0036) (0.0035)
Dep. var. mean 0.3538 0.3545 0.3830 0.0247 0.0276
Experienced Treated $\times$ Post -0.0191 -0.0164 -0.0144 -0.0045 -0.0025
(N = 141,184) (0.0222) (0.0224) (0.0231) (0.0053) (0.0049)
Dep. var. mean 0.3884 0.3919 0.4624 0.0454 0.0418
Pkg. contributor Treated $\times$ Post -0.0002 0.0033 0.0313 0.0108 -0.0166*
(N = 47,328) (0.0445) (0.0445) (0.0454) (0.0108) (0.0101)
Dep. var. mean 0.4802 0.4852 0.5865 0.0589 0.0556
@@
  • All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. files edited per PR and all intensive margin outcomes. Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023) Issue closed PR merge ratio
Overall Treated $\times$ Post -0.0010 -0.0014 -0.0052 -0.0016 -0.0070**
(N = 290,864) (0.0070) (0.0069) (0.0072) (0.0022) (0.0029)
Dep. var. mean 0.2164 0.2138 0.2444 0.0136 0.0347
Less experienced Treated $\times$ Post -0.0031 -0.0028 -0.0038 -0.0021 -0.0061*
(N = 149,680) (0.0097) (0.0096) (0.0099) (0.0023) (0.0033)
Dep. var. mean 0.2163 0.2144 0.2323 0.0083 0.0249
Experienced Treated $\times$ Post 0.0004 -0.0010 -0.0078 -0.0010 -0.0076
(N = 141,184) (0.0100) (0.0099) (0.0106) (0.0038) (0.0049)
Dep. var. mean 0.2164 0.2131 0.2573 0.0192 0.0452
Pkg. contributor Treated $\times$ Post 0.0096 0.0094 $0$0.0012 0.0020
(N = 47,328) (0.0189) (0.0185) (0.0198) (0.0082) (0.0099)
Dep. var. mean 0.2627 0.2582 0.3208 0.0271 0.0604
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post -0.0043 -0.0226 -0.0206 -0.0006 -0.0044
(N = 290,864) (0.0032) (0.0142) (0.0131) (0.0005) (0.0039)
Dep. var. mean 0.0381 0.1509 0.1383 0.0006 0.0606
Less experienced Treated $\times$ Post -0.0069* -0.0246 -0.0308* -0.0006 -0.0027
(N = 149,680) (0.0037) (0.0178) (0.0157) (0.0005) (0.0042)
Dep. var. mean 0.0300 0.1285 0.1087 0.0004 0.0315
Experienced Treated $\times$ Post -0.0015 -0.0213 -0.0079 -0.0007 -0.0062
(N = 141,184) (0.0053) (0.0228) (0.0217) (0.0009) (0.0069)
Dep. var. mean 0.0467 0.1747 0.1696 0.0009 0.0913
Pkg. contributor Treated $\times$ Post -0.0018 -0.0397 0.0249 -0.0021 0.0002
(N = 47,328) (0.0107) (0.0438) (0.0427) (0.0018) (0.0149)
Dep. var. mean 0.0656 0.2366 0.2193 0.0012 0.1376
@@
  • The placebo post-treatment period ranges from Mon 27 March 2023 to Thu 30 March 2023; the corresponding pre-treatment period is from Mon 20 March 2023 to Thu 23 March 2023. All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post 0.0109 0.0100 0.0083 -0.0008 0.0034
(N = 145,480) (0.0110) (0.0109) (0.0115) (0.0037) (0.0049)
Dep. var. mean 0.2492 0.2455 0.2872 0.0196 0.0424
Less experienced Treated $\times$ Post 0.0159 0.0152 0.0152 -0.0051 0.0054
(N = 58,400) (0.0174) (0.0174) (0.0179) (0.0045) (0.0061)
Dep. var. mean 0.2448 0.2424 0.2669 0.0114 0.0274
Experienced Treated $\times$ Post 0.0079 0.0069 0.0037 0.0024 0.0023
(N = 87,080) (0.0142) (0.0141) (0.0149) (0.0054) (0.0071)
Dep. var. mean 0.2522 0.2475 0.3008 0.0252 0.0524
Pkg. contributor Treated $\times$ Post 0.0326 0.0261 0.0225 0.0077 -0.0015
(N = 35,680) (0.0243) (0.0240) (0.0252) (0.0103) (0.0127)
Dep. var. mean 0.2853 0.2800 0.3456 0.0311 0.0609
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post 0.0021 0.0240 0.0413* $0$0.0071
(N = 145,480) (0.0050) (0.0232) (0.0214) (0.0005) (0.0066)
Dep. var. mean 0.0446 0.1718 0.1593 0.0008 0.0845
Less experienced Treated $\times$ Post 0.0028 0.0171 0.0354 -0.0003 -0.0037
(N = 58,400) (0.0067) (0.0338) (0.0271) (0.0005) (0.0081)
Dep. var. mean 0.0330 0.1382 0.1117 0.0006 0.0441
Experienced Treated $\times$ Post 0.0018 0.0297 0.0472 0.0002 -0.0095
(N = 87,080) (0.0070) (0.0315) (0.0310) (0.0008) (0.0096)
Dep. var. mean 0.0523 0.1943 0.1913 0.0010 0.1116
Pkg. contributor Treated $\times$ Post 0.0006 0.0170 0.0462 -0.0010 0.0037
(N = 35,680) (0.0128) (0.0572) (0.0493) (0.0012) (0.0178)
Dep. var. mean 0.0665 0.2376 0.2159 0.0014 0.1443
  • Notes:
  • The placebo post-treatment period ranges from Mon 11 April 2022 to Thu 14 April 2022; the corresponding pre-treatment period is from Mon 4 April 2022 to Thu 7 April 2022. All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
A: Any Activity
“Work” hours First activity Last activity Unusual activity (10th %tile) Unusual activity (25th %tile)
Overall Treated $\times$ Post 0.0356 -0.2048 -0.1933 0.0032 0.0053
(N = 290,864) (0.0497) (0.1777) (0.1784) (0.0038) (0.0063)
Dep. var. mean 0.9620 11.3925 14.5319 0.0494 0.1292
Less experienced Treated $\times$ Post 0.1344** -0.2833 -0.0351 0.0083 0.0130
(N = 149,680) (0.0643) (0.2543) (0.2506) (0.0052) (0.0084)
Dep. var. mean 0.8479 11.7579 14.6761 0.0477 0.1196
Experienced Treated $\times$ Post -0.0742 -0.1427 -0.3592 -0.0026 -0.0029
(N = 141,184) (0.0772) (0.2489) (0.2543) (0.0055) (0.0094)
Dep. var. mean 1.0829 11.0537 14.3982 0.0512 0.1395
Pkg. contributor Treated $\times$ Post 0.1684 -0.0484 -0.1477 0.0069 0.0226
(N = 47,328) (0.1535) (0.3616) (0.3987) (0.0107) (0.0180)
Dep. var. mean 1.4032 10.8529 14.4870 0.0584 0.1652
B: Productive Activities
“Work” hours First activity Last activity Unusual activity (10th %tile) Unusual activity (25th %tile)
Overall Treated $\times$ Post 0.0346 0.0159 0.0083 -0.0004 -0.0010
(N = 290,864) (0.0428) (0.2032) (0.2017) (0.0032) (0.0053)
Dep. var. mean 0.7395 11.4108 14.6531 0.0374 0.0955
Less experienced Treated $\times$ Post $0*$2090 -0.0059 0.0031 0.0082
(N = 149,680) (0.0567) (0.2848) (0.2782) (0.0045) (0.0073)
Dep. var. mean 0.6884 11.7987 14.7975 0.0383 0.0946
Experienced Treated $\times$ Post -0.0591 0.2347 0.0107 -0.0043 -0.0110
(N = 141,184) (0.0651) (0.2906) (0.2937) (0.0045) (0.0077)
Dep. var. mean 0.7937 11.0049 14.5019 0.0365 0.0965
Pkg. contributor Treated $\times$ Post 0.0822 0.0431 -0.2626 -0.0041 0.0075
(N = 47,328) (0.1296) (0.4233) (0.4664) (0.0086) (0.0150)
Dep. var. mean 0.9935 10.8366 14.4380 0.0416 0.1147
  • Notes:
  • All specifications include user–fixed effects, day-of-the-week–fixed effects, and a linear time trend for control and treatment group. “Productive activities” comprise: PullRequestEvent, PullRequestReviewEvent, PullRequestReviewCommentEvent, PushEvent, ReleaseEvent, CreateEvent, IssueEvent. The number of observations is depicted in parentheses after each sample definition. For first and last activity only users with at least one activity on the day are included; the number of observations for the “overall”, “less experienced”, “experienced”, and “pkg. contributor” sample in Panel A, respectively, Panel B are 43,868, 40,672, 84,540, 17,802 and 30,717, 32,135, 62,852, 12,663. Robust standard errors in parentheses are clustered on the user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Notes: Treatment effect estimates for the baseline DID specification are presented when users from one of the control group countries—i.e. Austria, France or Spain—are consecutively dropped from the sample. Point estimates and 95% (90%) cluster-robust confidence intervals are depicted in light (dark) grey for “Leave-One-Out” estimates and juxtaposed to the “Original” DID estimates (c. Table 1) in light (dark) red.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall (N = 290,864) Treated $\times$ Post 0.0062 0.0102 0.0090 -0.0045** 0.0045
[0.960] [0.602] [0.823] [0.083] [0.426]
Dep. var. mean 0.2314 0.2283 0.2624 0.0142 0.0359
Less experienced (N = 149,680) Treated $\times$ Post 0.0193* 0.0216** 0.0203* -0.0017 0.0118***
[0.179] [0.085] [0.157] [0.960] [0.000]
Dep. var. mean 0.2339 0.2316 0.2519 0.0089 0.0258
Experienced (N = 141,184) Treated $\times$ Post -0.0079 -0.0020 -0.0030 -0.0077* -0.0034
[0.960] [0.999] [0.998] [0.110] [0.971]
Dep. var. mean 0.2287 0.2249 0.2734 0.0198 0.0467
Pkg. contributor (N = 47,328) Treated $\times$ Post 0.0181 0.0164 0.0270 -0.0151* 0.0113
[0.945] [0.960] [0.710] [0.179] [0.835]
Dep. var. mean 0.2688 0.2644 0.3312 0.0267 0.0615
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall (N = 290,864) Treated $\times$ Post $0$0.0039 0.0220 -0.0004 0.0006
[0.999] [0.998] [0.311] [0.956] [0.999]
Dep. var. mean 0.0402 0.1581 0.1439 0.0007 0.0637
Less experienced (N = 149,680) Treated $\times$ Post 0.0040 0.0211 0.0456*** -0.0004 0.0037
[0.898] [0.823] [0.004] [0.823] [0.960]
Dep. var. mean 0.0315 0.1358 0.1148 0.0005 0.0340
Experienced (N = 141,184) Treated $\times$ Post -0.0046 -0.0324 -0.0031 -0.0003 -0.0028
[0.960] [0.581] [0.999] [0.996] [0.996]
Dep. var. mean 0.0494 0.1817 0.1749 0.0009 0.0953
Pkg. contributor (N = 47,328) Treated $\times$ Post -0.0221* -0.0933* 0.0226 -0.0021* 0.0127
[0.132] [0.109] [0.989] [0.126] [0.960]
Dep. var. mean 0.0661 0.2356 0.2210 0.0014 0.1393
  • Notes:
[…]
Appendix D GitHub Repository–User-Level Data